Momentum of an election is 600kev/c the group velocity and phase velocity of the electron is
1
Expert's answer
2020-10-12T07:44:10-0400

Solution

Energy of electron "E=0.511\times10^6eV"

Momentum of electron

"p=600\times10^3 eV/c"

Phase velocity is given by

"v_p=\frac{E}{p}=\frac{0.511\times10^6eV}{600\times10^3 eV/c}"

"v_p=0.255\times10^9m/s"

And

Group velocity is given by

"v_g=\frac{c^2}{v_p}=\frac{9\times10^{16}}{0.255\times10^9}"

"v_g=3.53\times10^8 m/s"